-0,044x³+2x²-50x+ax-600

Please solve for a and show calculations
1
Expert's answer
2011-03-23T05:53:39-0400
Let's try factorize this expression:
-0,044x³+2x²-50x+ax-600 = 2x²(-0,022x+1) +& (-50+a)x-600
Thus
-0.022 = (-50 + a)/(-600), -50 + a = 13.2, a = 63.2
Thus
2x²(-0,022x+1) +& 13.2x-600 = 2(-0,022x+1) - 600( - 0.022x +1) = 2 (x² -300)( - 0.022x +1)